# Chameleon Channels: Measuring YouTube Accounts Repurposed for Deception and Profit

### v1

@@ 3.1 The Fameswap Market
Verifying sales remains difficult, as with other scraping-based estimates [48, 49]. During our study, we scraped 16,110 seller profiles—7,000 more than Beluri et al. [7]—identifying 2,930 escrow transactions and 684 reviews. Only 1,590 transactions disclosed sale amounts, totaling USD 1.16M. Though modest, this likely represents only a fraction of actual sales, which also occur via forums, Discord, Telegram, and private deals on Fameswap.  [^1] We suspect public sales skew toward smaller, cheaper channels, as the average sale price (USD 731) is far below the average listing price (USD 5,400).
3.1.1 Account Ownership Verification
Fameswap verifies account ownership by providing sellers with a unique randomized string that should be placed in the advertised social media profile. For example, a seller receives the string aSgc5H3s and places it temporarily in their YouTube channel description, allowing Fameswap to verify the ownership. This process is similar to using TXT records to verify domains.
3.1.2 Other Social Media Account Markets
Chu et al.[8] identified five marketplaces for social media accounts: SWAPD[50], Accs-Market.com [51], Trustiu (now inactive), ViralAccounts [52], and Fameswap [53]. We compare primarily with Chu et al.’s work, as Beluri et al.’s study became public after our data collection began and is referenced retrospectively. We also identified additional venues, including OGUser (similar to SWAPD) and Telegram channels.
Of these, only Fameswap and Accs-Market offer e-commerce-style interfaces, with listings, escrow, and reputation systems. ViralAccounts operates as a broker, and SWAPD functions as a private forum, limiting large-scale analysis. At study onset (October 2024), Accs-Market was inaccessible (HTTP 500), and archived pages returned HTTP 301 errors, preventing further analysis. We therefore focused on Fameswap, given its accessibility, scale, and apparent prominence.
3.2 YouTube Handles and IDs
In October, 2022, YouTube introduced handles as a way to uniquely identify accounts [54]. Different from channel titles, handles (“@SomeChannel”) are unique and can be accessed through “www.youtube.com/@SomeChannel.” Handles were rolled out gradually, but by the end of 2023 most accounts would have chosen or received a unique handle for their account. Handles resolve to a unique channel ID, allowing a channel to change its handle but keep pointing to the same channel. YouTube’s channel ID is a string that begins with “UC”, followed by 22 characters (letters, numbers, dashes, and underscores). Channel IDs can also be used to access a channel in the form: “www.youtube.com/channel/UC…”. Importantly, while an account can change its handle, its channel ID cannot change, which allows us to monitor accounts over time.
4 Methods
We measure the repurposing of YouTube accounts and their participation in disseminating problematic content through the lens of sold accounts extracted from a social media marketplace (Fameswap) and by observing handle changes and channel repurposing in the wild (from Social Blade). Specifically, we structure our study in three parts. First, we define, characterize, and estimate the prevalence of channel repurposing (Section 5). Second, we qualitatively annotate channels to identify potentially problematic content based on YouTube’s guidelines and prior work (Section 6). Third, we describe our framework for deriving statistically valid estimates using textual annotations and quantitatively test what features are indicative of repurposing (Section 7). An end-to-end view is later provided in Figure 9.
This section discusses our data collection practices and each sample used in the study. The following sections describe each experiment in turn.
4.1 Data Collection and Sources
4.1.1 YouTube Channels For Sale on Fameswap
As illustrated in Figure 2(a), we conducted daily scrapes of Fameswap from October 26, 2024, to March 31, 2025. The Fameswap interface provides a paginated list of all account listings. We scraped all historical listings. This set includes all listings not deleted or hidden from the website. On October 2024, Fameswap began displaying channel IDs with YouTube listings (whereas before they only disclosed the channel title). Together with our Fameswap scrapes, we began conducting regular scrapes of all YouTube channels advertised across Fameswap listings. We scraped channels every 3 days on average. As of March 31, 2025, we collected 4,641 YouTube channel IDs advertised for sale, each with about eight observations per month. For each channel we collected from Fameswap, we scraped their YouTube channel using the YouTube Data API v3.
4.1.2 Collecting a Large Sample of YouTube through Social Blade
To estimate the prevalence of channel repurposing and identify repurposing indicators beyond marketplaces, we aimed to observe this phenomenon “in the wild.” Randomly sampling YouTube is difficult due to the lack of a centralized channel directory and method to enumerate channel IDs [55]. Horta Ribeiro and West address this by sampling from Social Blade [56], an analytics site that has indexed over 68M channels in the past 17 years [57]. Social Blade is a widely used reference for creators [58] and has informed multiple YouTube studies [59, 60]. Building on this method, we sampled a large set of channels from Social Blade using a sample from Horta Ribeiro and West. While Social Blade’s crawling process is not public, it likely favors more popular channels. This, however, is acceptable given that our focus is content creators with larger audiences.
We sampled 1.4M channels and conducted an initial scrape from December 23, 2024, to January 21, 2025 (Figure 2(b)). We collected metadata from 1,397,586 channels and 139M corresponding videos. A second scrape, from March 21–31, 2025, captured updated metadata and 20M new videos. Only 1,351,912 channels returned data, indicating a 3.3% deletion rate. We used yt-dlp, a common tool for archiving YouTube content [61, 62, 63], discussed further in Section 9.
4.1.3 Samples Considered in this Study
  • Set of channels for sale on Fameswap: All channels collected from Fameswap, $n$=4,461 from October 21 2024, to March 31 2025 (see Section 4.1.1).
  • Set of repurposed channels from Fameswap: Subset of repurposed Fameswap channels, $n$=1,084 (see Section 5.2.1).

### v2

@@ 5 Fameswap Results
Verifying sales remains difficult, as with other scraping-based estimates [17, 29]. We scraped 16,110 seller profiles, 7,000 more than Beluri et al.—like them, we also observed inflated pricing for some listings: 41 listings exceeded USD 100K, and 16 exceeded USD 1M [12]. We identified 2,930 escrow transactions and 684 reviews. Only 1,590 transactions disclosed sale amounts, totaling USD 1.16M. Though modest, this likely represents only a fraction of actual sales, which also occur via forums, Discord, Telegram, and private deals on Fameswap  [^3] We suspect public sales skew toward smaller, cheaper channels, as the average sale price (USD 731) is far below the average listing price (USD 5,400)..
5.1 Before and After Repurposing
Using our repurposing definition (§ 4.2) and our annotations (§ 4.4), we labeled the set of Fameswap channels we collected. We identified 1,024 Fameswap channels (23%) that were repurposed during our observation period. In total, these channels had a combined audience of 193M+ subscribers. In the following, we explore how long it took for these channels to change, the impact of repurposing on their existing audience, whether they were suspended, and the content to which they shifted.
Figure 5: Days since a channel is listed for sale, until it changes its handle, title, and description (cumulative).
Time Until Repurpose. For each listed channel, we computed the time it took them to change their handle, title, and description. As seen in Figure 5, the median time for a handle change was 28.5 days, followed by 29.3 days for a change in the handle and title, and 30 days for a change in handle, title, and description. Interestingly, handle, title, and description changes are not instantaneous, but instead occur over a period of 36 hours (1.5 days), a fact that could inform an anomaly detection system. These results can be interpreted as a proxy for the median time of a sale.
Figure 6: Subscriber growth comparison between repurposed (treatment) and non-repurposed channels (control). X-axis is relative to repurposing event. For non-repurposed accounts, the X-axis is based on the time they were first listed on Fameswap. Percentage growth is relative to the number of subscribers at tt=0.
Impact of Repurposing on Subscribers. To assess the impact of repurposing on subscriber count, we aligned all time series so that $t=0$ marks the week of the channel’s change (Figure 6). We tracked subscriber change (%) over the four weeks before and 12 weeks after this point. We compared the subscriber growth against a control group composed by channels from Fameswap that were not repurposed. For these channels, we defined tt=-4 as the first week when listed on Fameswap and we observed them for a total of 16 weeks (up to tt=12).
Most channels show slight but steady growth prior to the change, likely due to residual recommendation rather than active content updates. This pattern persists in the control group. However, repurposed channels’ subscriber counts increase on average–—often substantially. While some unsubscribes may be masked by new subscribers, the consistent growth suggests most users remain unaware of the change and stay subscribed. Lastly, on average we do not observe sharp changes in subscriber counts, which have been associated with the purchase of artificial engagement services (e.g., buying bot followers in bulk [60]), or a large decrease due to a bulk suspension of bot accounts.
Rate of Suspension. We estimate the rate at which YouTube suspends Fameswap accounts, repurposed or otherwise. To do so, we compute the survival function for two groups, repurposed channels and those listed but not repurposed during our observation period, using the Kaplan-Meier estimator. As shown in Figure 7, all Fameswap channels exhibit an average survivability of approximately 85% after 200 days (95% CI: 79%–87.5%). Based on a log-rank test, there is no significant difference between repurposed channels and those that have not yet been repurposed. Given that the unofficial sale of accounts likely violates YouTube’s guidelines, we expected a higher rate of suspensions. However, these results indicate that repurposed channels are currently avoiding penalties.
Figure 7: Survivability curve for repurposed and non-repurposed Fameswap channels. “Death” event is channel removal. Bands represent 95% CI.
Figure 8: Topic presence for reused Fameswap channels. Topics on the left are self-reported at the time of listing. Topics on the right are labels assigned during our topic detection. Flows are proportional and one-to-many. That is, if a channel has several topics, flows are drawn to all destination topics proportionally. Counts overlap because some channels may have presence in both categories.
Changes in Content. We trace how channels transition from their original content category to policy-sensitive content after being repurposed. Using self-reported Fameswap categories (e.g., “gaming”, “technology”, etc.) and annotations after a channel has been repurposed, we highlight how content shifts.
When listing a channel for sale on Fameswap, users choose a category that best fits the channel’s content. Choosing a category improves the searchability of the channel in the marketplace. As observed in Figure 4, there are 18 categories that cover a wide range of topics. More than 88% of the listings have categories. We manually verified the validity of these self-assigned categories and found them to be largely accurate. Using these self-assigned categories, we investigated transitions in content categories, as shown in Figure 8. We find that 545 (53%) of repurposed channels later displayed policy-sensitive content in their channels. Of these, 374 had ideological content and 286 had financial content (non-exclusive). In particular, the origin category (from Fameswap) had little to do with the types of content that would appear in the channel. We see that every category, even though seemingly benign (e.g., “motivation”, “food”) can shift to posting policy-sensitive content.
Not all channels began posting policy-sensitive content. During our observation period, there were many channels in which we detected no policy-sensitive content even after being repurposed (n=n=479). For example, a sports page became a musician’s page, and an entertainment page later became an influencer’s travel vlog. Some repurposed channels would later become channels associated with personal or commercial brands. Although not posting policy-sensitive content, these transitions raise a question about transparency across channels with large followings, given that people may rely on metrics such as subscriber counts to judge the authenticity of a brand.
6 Generalizability and Prevalence
Table 1: Regression results and descriptives for Fameswap and Socialblade models.
Fameswap Socialblade Fameswap (n=1,024) Socialblade (n=1,047) Baseline (n=2,972)
Coefficient (SE) Coefficient (SE) Mean SD Mean SD Mean SD
(Intercept) $\phantom{-}0.7093^{***}$ $(0.0994)$ $\phantom{-}0.3401^{***}$ $(0.0824)$
View Ct. (Mean) $\phantom{-}0.0000\phantom{{}^{***}}$ $(0.0000)$ $\phantom{-}0.0000\phantom{{}^{***}}$ $(0.0000)$ 114,474 615,572 119,214 1,016,701 211,258 1,085,125
View Ct. (SD) $\phantom{-}0.0000\phantom{{}^{***}}$ $(0.0000)$ $\phantom{-}0.0000\phantom{{}^{***}}$ $(0.0000)$ 379,984 1,298,554 195,928 1,100,404 442,672 2,061,892
Like Ct. (Mean) $\phantom{-}0.0000^{***}$ $(0.0000)$ $\phantom{-}0.0000\phantom{{}^{***}}$ $(0.0000)$ 4,085 19,742 1,120 8,357 1,657 9,986
Like Ct. (SD) $\phantom{-}0.0000^{***}$ $(0.0000)$ $\phantom{-}0.0000\phantom{{}^{***}}$ $(0.0000)$ 12,994 53,193 1,761 9,201 3,290 15,435
Comm. Ct. (Mean) $0.0007^{*}\phantom{**}$ $(0.0005)$ $0.0005\phantom{{}^{***}}$ $(0.0004)$ 54.78 301.1 45.09 270.5 88.28 400.0
Comm. Ct. (SD) $\phantom{-}0.0001^{*}\phantom{**}$ $(0.0002)$ $\phantom{-}0.0000\phantom{{}^{***}}$ $(0.0002)$ 180.41 982.0 73.72 303.6 166.17 752.0
Vid. Len. (Mean) $\phantom{-}0.0001^{*}\phantom{**}$ $(0.0000)$ $\phantom{-}0.0000\phantom{{}^{***}}$ $(0.0000)$ 739.57 2,821.4 518.60 1,207.4 616.16 1,113.2
Vid. Len. (SD) $\phantom{-}0.0000^{***}$ $(0.0000)$ $\phantom{-}0.0000\phantom{{}^{***}}$ $(0.0000)$ 1,316.0 12,617.2 580.82 1,489.1 803.19 8,011.8
T. Btw. Vids. (Mean) $0.0039^{*}\phantom{**}$ $(0.0022)$ $\phantom{-}0.0002\phantom{{}^{***}}$ $(0.0002)$ 17.59 58.93 119.73 328.23 72.11 213.58
T. Btw. Vids. (SD) $\phantom{-}0.0050^{***}$ $(0.0013)$ $\phantom{-}0.0022^{***}$ $(0.0003)$ 41.64 108.8 206.89 381.47 115.11 200.46
Vids./Wk. (Mean) $\phantom{-}0.0046\phantom{{}^{***}}$ $(0.0055)$ $\phantom{-}0.0193^{*}\phantom{**}$ $(0.0117)$ 5.65 9.03 3.01 5.59 3.65 18.43
Vids./Wk. (SD) $\phantom{-}0.0058\phantom{{}^{***}}$ $(0.0048)$ $0.0047\phantom{{}^{***}}$ $(0.0053)$ 4.98 10.20 2.37 6.22 3.08 20.60
T. 1st to Oldest Vid. $\phantom{-}0.0289^{***}$ $(0.0050)$ $\phantom{-}0.0224^{***}$ $(0.0027)$ 17.78 23.15 57.72 64.77 84.15 62.78
Oldest Vid. Age $0.0744^{***}$ $(0.0045)$ $0.0375^{***}$ $(0.0025)$ 20.10 23.43 65.78 68.18 132.43 56.53
Subs. Count $\phantom{-}0.0000\phantom{{}^{***}}$ $(0.0000)$ $\phantom{-}0.0000^{*}\phantom{**}$ $(0.0000)$ 188,898 1,016,056 34,056 169,278 103,427 1,002,724
Total Videos $0.0001^{*}\phantom{**}$ $(0.0001)$ $0.0002\phantom{{}^{***}}$ $(0.0001)$ 162.76 493.97 127.09 806.19 373.66 1,365.7
Content Features
Non-YT Money $\phantom{-}0.5092^{***}$ $(0.1464)$ $\phantom{-}0.0713\phantom{{}^{***}}$ $(0.0936)$ 0.347 0.476 0.372 0.483 0.432 0.495
AI-generated $\phantom{-}1.1557^{***}$ $(0.2087)$ $\phantom{-}0.6283^{***}$ $(0.2000)$ 0.112 0.316 0.066 0.248 0.022 0.147
Political $\phantom{-}0.1697\phantom{{}^{***}}$ $(0.1799)$ $0.1145\phantom{{}^{***}}$ $(0.1303)$ 0.217 0.412 0.171 0.377 0.187 0.390
Religious $\phantom{-}0.9681^{***}$ $(0.1645)$ $\phantom{-}0.3910^{***}$ $(0.1187)$ 0.208 0.406 0.194 0.396 0.154 0.361
News $\phantom{-}0.7261^{***}$ $(0.2001)$ $\phantom{-}0.4305^{***}$ $(0.1356)$ 0.183 0.387 0.154 0.361 0.129 0.335
Medical/Health $0.4594^{**}\phantom{*}$ $(0.1921)$ $0.3025^{*}\phantom{*}$ $(0.1300)$ 0.096 0.294 0.121 0.327 0.153 0.360
Cryptocurrency $\phantom{-}1.7541^{***}$ $(0.2622)$ $\phantom{-}0.5112^{*}\phantom{**}$ $(0.2424)$ 0.116 0.321 0.039 0.194 0.014 0.118
Gambling $\phantom{-}1.4821^{***}$ $(0.3274)$ $\phantom{-}0.2759\phantom{{}^{***}}$ $(0.2728)$ 0.056 0.229 0.021 0.143 0.017 0.129
Money/Stocks $\phantom{-}0.2390\phantom{{}^{***}}$ $(0.1892)$ $\phantom{-}0.1741\phantom{{}^{***}}$ $(0.1451)$ 0.162 0.369 0.113 0.316 0.073 0.260
Kids $\phantom{-}0.6960^{**}$ $(0.2275)$ $\phantom{-}0.3357^{**}\phantom{*}$ $(0.1341)$ 0.099 0.298 0.1242 0.3299 0.075 0.264
P. Copyright Infr. $\phantom{-}0.5022^{***}$ $(0.1517)$ $\phantom{-}0.2826^{**}\phantom{*}$ $(0.0943)$ 0.254 0.436 0.345 0.476 0.426 0.495
Manosphere $\phantom{-}0.8533^{*}\phantom{**}$ $(0.4041)$ $\phantom{-}0.4109\phantom{{}^{***}}$ $(0.4444)$ 0.036 0.187 0.011 0.106 0.006 0.078
  • Notes: $^{***}p<0.001$, $^{**}p<0.01$, $^{*}p<0.05$ ; Overlapping channels across samples were excluded.
We investigate the prevalence of channel repurposing and whether channel repurposing is observable beyond the accounts that are traded in social media marketplaces like Fameswap. By using a random sample of channels from the analytics platform Social Blade, we estimate the prevalence of repurposing “in-the-wild.” Furthermore, using a logistic regression model we investigate which channel features correlate with repurposing. We confirm that repurposed accounts, beyond those that are commercialized in open markets, are also used to disseminate the content we identified in § 4.3. In addition, we find that repurposed channels tend to have more subscribers on average. Lastly, posting behavior (e.g., upload cadence) and content are most predictive of repurposing. We summarize the descriptive statistics of our three samples (repurposed Fameswap channels, repurposed Social Blade channels, and a random sample of non-repurposed Social Blade channels) in Table 1.
6.1 Prevalence of Repurposed Channels in the Wild
We labeled the content of the large sample of YouTube channels we collected in § 4.1.2, using the annotation procedure in § 4.2. Of the 1.4M YouTube channels, we detected 10,200 (0.73%) channels that changed their handle between January and March 2025. Of these, our LLM labeled 3,584 channels as repurposed. To validate the annotations, we randomly sampled 10% of the channels (nn=1,020) and manually annotated each pair of observations according to our definition. The false positive and negative rates are 4.9% and 3.8%, respectively.
The agreement between our annotations and the LLM classifier is 95.54%. Our classifier had an accuracy between 94.2% and 96.6% (95% CI), resulting in an estimated 3,384–3,456 repurposed accounts. These results indicate that, from Jan. 2025 to Mar. 2025, approximately 0.24-0.25% of channels in the Social Blade population (68M+ channels) were repurposed. Of the repurposed channels we observed, 1,074 had more than 1,000 followers. These repurposed channels had a total audience of 43,975,420 subscribers, a sizable audience. These results imply that there were 160,000+ repurposed channels ($\sim$0.24% of the 68M captured by Social Blade) that we did not capture, likely with tens or hundreds of millions of subscribers.
6.2 Indicators of Repurposing
We use a logistic regression model to examine the relationship between channel characteristics, channel content, and channel repurposing. We model the log-odds of the outcome as a linear combination of the predictors:
where $\beta_{0}$ is the intercept (log-odds outcome when all predictors are zero), $F_{i}$ represents quantitative features extracted from each channel, and $T_{i}$ represents an indicator for each content category. In this model, a one-unit change in a covariate $X_{i}$ changes the log-odds outcome by $\beta_{i}$; equivalently, the odds change by a factor of $e^{\beta_{i}}$.
Quantitative Features. We take the mean and standard deviation of all observed video views, likes, and comments. We also compute the mean and standard deviation of the number of videos posted weekly, video lengths (seconds), and the time between each video posted (days). Finally, we compute the time between the first and last video the channel has posted (months) and the oldest video the channel has, which would typically mark when the channel first started uploading content. Finally, we take the total number of videos. Table 1 presents the descriptive statistics for each sample. Video deletions are likely an important feature. However, because we do not have frequent snapshots of Social Blade channels (given the sample size), we miss all uploads and deletions that take place between the two snapshots—a limitation that the time between videos may still partially address.
Qualitative Features. For content, we encode topic presence as indicator variables derived from our LLM categorization and gold-standard labels (described in § 4.3).
6.3 Regression Results
We conduct two logistic regressions, first we model repurposed channels from Fameswap against a baseline drawn from Social Blade using the R package,  DSL [21]. We report the coefficients in Table 1.
Takeaways from Quantitative Features: Channels whose oldest video is more recent are more likely to be repurposed, as are those with a longer interval between the oldest and most recent uploads. So, buyers seemingly favor channels that look established (i.e., showing a sizeable upload time span), but those are not necessarily truly long-standing: a genuinely mature channel would feature both an old first upload and a long production span.
Irregular upload intervals (high SD in “time-between-videos”) predict repurposing, implying that erratic schedules and an existing audience are typical markers of repurposed channels. Average views across videos were not significantly different across samples. Repurposed Fameswap channels tended to have a higher number of likes in videos, a greater variance in likes in videos, and lower comment counts. It is possible that repurposed channels may disable or limit comments to avoid scrutiny. The likes and views pattern did not hold for repurposed channels found in the wild. However, in both cases subscriber counts were not a predictor of channel reuse, indicating that channels of all sizes are repurposed. However, sold channels tended to have higher subscriber counts possibly because they are more appealing for sale.
Takeaways from Qualitative Features: Repurpose odds are primarily content-driven. AI-generated content ($\beta$=0.6–1.1) is the strongest predictor of repurposing, probably because automated pipelines scale cheaply after a channel is acquired. However, we rely on AI usage disclosure to tag channels, as described in § 4.3, so the coefficient may also reflect self-disclosure bias. Cryptocurrency shows the next-largest effect ($\beta$=1.7 on Fameswap), followed by gambling. We hypothesize that these categories are prone to reuse because of their monetization potential. In addition, dubious projects may repurpose accounts to be perceived as legitimate. In line with previous work, we find a significant amount of off-platform monetization [40, 15], albeit only for Fameswap. Medical/health is negatively associated, suggesting either stricter platform scrutiny or lack of prominent health-related narratives. Religious content and potentially copyright-infringing content are associated with repurposing. A possible explanation could be that these types of content are used to grow channels due to their appeal (e.g., free TV shows), similar to kids content, and we are identifying trace videos prior to deletion.
7 Discussion and Conclusions
Our results suggest that channel repurposing involves sold channels (such as those advertised on Fameswap), as well as channels that may not have been for sale (such as most of those found through Social Blade). Across most Fameswap accounts, channel repurposing seemed to go unnoticed both by users and YouTube, as indicated by the subsequent growth in subscribers and lack of suspension. The median time between a channel being listed for sale and being repurposed was 30 days. However, some took as long as 200+ days (RQ1).
In both samples, we find that channels with large numbers of subscribers—193M+ for Fameswap and 43M+ for Social Blade—completely repurposed their channels, erasing any perceivable association with their prior identity. By leveraging a large sample of YouTube channels, from a random Social Blade sample, we estimated that 0.24-0.25% of channels were repurposed between January and March 2025. These results imply that there were another 160,000+ repurposed channels (out of the 68M captured by Social Blade) that we did not capture, likely with tens or hundreds of millions of subscribers (RQ2).
Engagement metrics offer little indication that a channel has been or will be repurposed. Instead, video upload behaviors (i.e., time between videos) and potential gaps in their video history are better indicators for potential repurposing. These indicators, unfortunately, may not be detectable or readily interpretable by everyday users. Furthermore, usage of AI was also a significant predictor of channel repurposing, indicating that “faceless channels” are commonly repurposed (RQ3).
The presence of policy-sensitive content in the channel was also a significant predictor in both samples. On Fameswap, 53% of channels posted potentially deceptive ideological and financial content, as per YouTube’s guidelines. Before repurposing, they belonged to various innocuous content categories, ranging from humorous content to sports and celebrity gossip. Figure 1 illustrates an example of a problematic transition, in which a channel that previously shared “interesting facts” is sold and repurposed into a news channel featuring contentious geopolitical actors: Russia, Ukraine, Zelenskyy, etc. Due to space constraints, we cannot discuss all remarkable cases individually, and instead, quantitatively show that policy-sensitive content is significantly related to channel repurposing (H1).
Our results suggest that social media audiences are becoming more commoditized (i.e., have less differentiation, have widespread availability, and compete mainly on price). From the lens of transaction cost economics, social media account markets are evolving from forums (unassisted markets) to assisted markets, with lower barriers to entry and increased liquidity [76]. This transition typically makes outsourcing more attractive and, as a result, fosters specialization, a pattern that frequently emerges in cybercriminal endeavors [81, 72, 41]. In our context, this may signal the emergence of actors who specialize in harvesting organic audiences.
We do not claim nor expect these markets to be free from fraud. Many vendors may use artificial engagement to inflate their numbers and sell accounts at a premium. However, these markets are growing. They are also increasingly becoming more sophisticated, allowing sellers to share screenshots and incorporating feedback systems. This continuous evolution, coupled with the apparent success of these markets, seems to indicate that a substantial portion of buyers find these markets useful.
We found evidence that repurposed channels may play a role in influence operations, which, similar to past work, is driven by content features [8]. We confirm this finding in the context of YouTube by uncovering various channels that turned into journalistic-like channels. Additionally, generative artificial intelligence (GenAI) seems to play an important role in the automation of content creation to harvest audiences, in line with recent small-scale investigations [19]. Putting misinformation consumption in terms of supply and demand, scholars have argued that AI does not increase the demand for misinformation [68]. However, our results add nuance to these claims. By facilitating the cultivation of audiences that can then be served mis/disinformation, GenAI may be helping increase the demand for disinformation indirectly. Furthermore, GenAI may enable actors to more easily produce diverse content across niches and languages, facilitating the cultivation of specific audience demographics—a strategy that can be employed to target specific populations when selecting channels for repurposing.
There are a few possible interventions. First, our results highlight a set of features that distinguish repurposed channels from benign ones. Second, we recommend increased transparency. There are many legitimate reasons to change channel handles and titles, and doing so without drawing attention is somtimes advisable. However, we show that current imperceptibility can be exploited to the detriment of users. Thus, if a channel is large enough, we suggest that subscribers be notified of meaningful changes in identity and content, and that historical handle changes and titles be auditable. Another option is to monitor second-hand social media account marketplaces and seek to deplatform them. We do not suggest this approach. These interventions often have short-lived benefits and make nefarious activity harder to measure [84]. Lastly, another option is to make monetization harder, a strategy that YouTube adopted on July 15th, 2025, to target inauthentic content [31]. However, many accounts already leverage various forms of off-platform monetization. Out of these interventions, we are most optimistic about approaches that increase transparency and user agency.
7.1 Limitations and Future Work
We rely on video metadata to infer topics, without analyzing the video content itself. Thus, when labeling channels as political, medical, or news-related, we do not assess whether the content is actually harmful or mis/disinformation. Automated fact-checking is a hard and open problem [93], particularly for developing events. In addition, fact-checking video content remains underexplored, with most of the work focused on news articles [35]. Nonetheless, the veracity of the claims does not preclude the usefulness of these channels as vectors of influence.
We only checked for account suspension as a punitive measure. However, de-monetization likely precedes suspension. To the best of our knowledge, there is currently no reliable method to check monetization status, which inhibited us from verifying whether YouTube is currently detecting repurposed channels and de-monetizing them. However, even if this is the case, account marketplaces have only gotten more popular, which puts into question whether de-monetization is a sufficient deterrent.
Assessing harm is subjective and verifying whether exposure to content led to harm is often impossible. While not all cryptocurrency or money-themed channels promoted obvious scams, past work has financial loss back to influencers’ advice [45]. To assess claims in videos, future work could analyze transcripts, video frames, or external URLs to refine these labels and estimate the presence, volume, and degree of harmful content.
Fameswap seems to be among the leading marketplaces [12]. However, the extent of overlap across marketplaces or between marketplaces and forums is unknown. This motivated our prevalence estimation, which avoids limiting analysis to marketplace samples. Finally, our study focuses only on YouTube. While we expect our findings to generalize to platforms like TikTok, Twitch, or Twitter/X, this remains to be tested.
Acknowledgments. This research was partially supported by the Singapore Defence Science and Technology Agency (DSTA) under agreement CNZ2000832. It benefited from great work by Haoxiang Yu. Lastly, immense thanks to our external social media expert, Asia Grant.
Appendix A Ethical Considerations
This study involved various stakeholders and careful ethical consideration of various decisions, particularly data collection. Ultimately, our decision to conduct and publish this study is guided by the fact that it is the first to document repurposing of social media accounts, which we find may affect millions of people. All data collected were stored on a private server hosted at our university and accessed only by members of the research team, unless otherwise indicated.
Fameswap. Fameswap is a website allegedly registered in the US with 12 years of activity. Purchasing a Fameswap membership raises ethical concerns, as it supports a service that some individuals can use to carry out policy-violating practices, as we found in this study. However, Fameswap is not illegal nor do we claim that Fameswap is an active collaborator in the practices we uncovered. However, as long as 1) alternative options are not available, and 2) the monetary amounts considered are small—especially when compared to the potential scientific value of developing countermeasures—there is precedent in the research community, even for goods and services that are illegal, such as “booters” (DDoS-for-hire services) [44], goods from spam-advertised websites [49, 43], and even hacking services [58], with purchases surpassing a thousand dollars. These numbers are in line with or exceed our own expenses: USD 348 for a monthly and yearly subscription.
The Fameswap platform does not explicitly prohibit scraping and there is also recent precedent of account registration to access data, even in problematic platforms [13, 37, 30]. The site’s robots.txt disallows nothing (i.e., Disallow: <empty>). Our web scraping activities added to server load and bandwidth usage; however, we collected at most tens of pages per day; we believe our impact was negligible.
We considered whether to publish Fameswap’s name, as it may attract attention to it. We decided to name Fameswap as it has been named in recent work [15, 12] and may encourage researchers to study similar platforms and other commercialized accounts beyond YouTube. Distribution of the data we collected from Fameswap, however, merits scrutiny. We expand on this point in § B.
Users of the Fameswap Marketplace. In this study, we collected metadata from YouTube accounts listed for sale on Fameswap, as well as metadata from those accounts through YouTube. We did not attempt to deanonymize any Fameswap user nor did we employ user-level data beyond counting the number of vendors on the platform. We did not interact with users (e.g., via direct messages). We do not report quotes. We blur account and channel identifiers, where necessary, to prevent driving traffic to these channels and for user privacy.
Social Blade and the Internet Archive. Similar to Fameswap, we sought to mitigate our impact on the platforms from which we collected data: the Wayback Machine and Social Blade. Where possible, we opted to use existing data, rather than query it ourselves. We did not scrape Social Blade, but instead sampled from the data that Horta Ribeiro and West had collected [39]. To collect data from the Internet Archive, we followed their API guidance and limits [1].
YouTube. We employed yt-dlp to assist in obtaining data from YouTube during this period. It is important to note that historically there has been some concern over the use of yt-dlp. In 2020, the Recording Industry Association of America, issued a takedown notice to GitHub under the Digital Millenium Copyright Act (DMCA), requesting the removal of the project and its forks, arguing that it violated German copyright law [91]. Nonetheless, the takedown was reversed [91], and is worth noting that the main concern was using yt-dlp for video downloads, not metadata (which was our use). To date, yt-dlp remains a useful tool for academic projects, even for video downloads [64]. In our case, we only used yt-dlp to collect metadata from channels and chose conservative limits. We originally intended to collect data through the YouTube Research program. We submitted a proposal prior to starting the study and received approval. Unfortunately, there was an issue with the API key that was provided and this was only resolved after the study period ended.
YouTube Users. For all YouTube channel data, we sought to incorporate the privacy guidelines and considerations proposed by Beadle et al. [11]. Although YouTube does not meet Beadle et al.’s definition of social media data[11, 88], we still applied the same privacy considerations as with other social media sites. We did not use emails or phone numbers in our analysis, even though some accounts volunteer this information in their channel metadata. We also focused our analyses on larger channels (>>1,000 subscribers), as we believe that they have a lower expectation of privacy, given their large audience.
Research Team. Labeling the content in each channel required data annotation which was conducted primarily by two members of the research team and one external member. All annotators were briefed about the potential content they would encounter. Each annotator participated voluntarily, eagerly, at their leisure, and with the freedom to stop at any point without repercussions. Because we only focused on metadata, the chances of encountering disturbing content were minimized.
Appendix B Open Science
To conduct this study, we pre-registered our regression hypotheses, methods, and analysis plan prior to conducting the study. The pre-registration is available at https://doi.org/10.17605/OSF.IO/FG9PM. Our Zenodo repository with supplemental material is available at https://doi.org/10.5281/zenodo.18216091.
B.1 Samples Considered in this Study
The main samples that we considered in this study are:
  • Set of channels for sale on Fameswap: All channels collected from Fameswap, nn=4,641 from October 21 2024, to March 31 2025 (see § 4.1.1).
  • Set of repurposed channels from Fameswap: Subset of repurposed Fameswap channels, nn=1,024 (see § 4.2).